OpenStudy (jordan238):

one number is 4 less than 3 times another. their sum is -16. find the smaller

OpenStudy (kayleerose0715):

are you supposed to write an expression?

OpenStudy (kayleerose0715):

3(n-4)=-16 n will equal -1.3 repeating

OpenStudy (radar):

Let x =the another number.........then 3x is three times the another number........ then 3x -4 is the one number.......now we can say 3x-4 + x is the sum of one number and another number The problem tells us that sum is -16 3x - 4 + x = -16 4x - 4 = -16 add 4 to both sides of equal sign. 4x = -12 divide both sides by 4 x = -3, "another" number is -3 -9 - 4 = -13 which will be the "one" number (3x - 4)